# Revisiting Stochastic Approximation and Stochastic Gradient Descent

### v2

@@ 1.1 Historical Review
$\hat{G}(Y_{n})=\frac{1}{2c_{n}}[(F(Y_{n}+c_{n})+M^{\prime}_{n+1})-(F(Y_{n}=c_{n})+M^{\prime\prime}_{n+1},$ (3)
@@
Subsequently, Blum [3, 4] extended the approach to the of arbitrary $d$, and also proved that the iterations converge almost surely, provided (5) hold. For this reason, the conditions (5) are referred to as the Kiefer-Wolfowitz-Blum (KWB) conditions.
@@ 1.2 Contributions of the Paper
    In the special case where $\{W_{n}\}$ are i.i.d.  with a symmetric distribution (that is, $W_{1}$ and $-W_{1}$ have the same distribution), GSLLN holds if there exists a $\delta\in(0,1]$ such that
@@ 2.2 Stochastic Approximation - Main Result
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Let $x^{*}$ denote the unique fixed point of $H$, and define $G(x)=x-H(x)$. Then clearly $x^{*}$ is the unique solution of $G(x)={\bf 0}$. Moreover, it is easy to see that (11) is satisfied with $b=c=1$. This is because
@@
where for symmetric matrices $A$ and $B$, $A\leq B$ is equivalent to $B-A$ being positive semidefinite. The above condition is also equivalent to the statement that the largest eigenvalue of $M(x)$ is bounded above by $h<\infty$, and the smallest eigenvalue of $M(x)$ is bounded below by $g>0$. With these assumptions, it is shown that (7) holds with $\|\cdot\|=\|\cdot\|_{2}$, and
$b=\frac{1}{1+g+h},\quad\rho=\frac{g}{1+g+h}<1.$
@@
[…]
@@
[…]
By assumption, the spectrum of the symmetric matrix $M(x^{*}+\lambda(x-x^{*}))$ lies in the interval $[g,h]$. Hence the spectrum of $I-bM(x^{*}+\lambda(x-x^{*}))$ lies in the interval $[1-\frac{h}{1+g+h},1-\frac{g}{1+g+h}]$. Since both numbers are positive, the spectral norm (which is the same as the $\ell_{2}$-induced norm) satisfies
$\|I-\rho M(x^{*}+\lambda(x-x^{*}))\|_{2\rightarrow 2}\leq 1-\frac{g}{1+g+h}=\rho.$
@@
[…]
To motivate the concept of the GSSLN, let us study a special case, when $G(x)=tx$, $t>0$ and $\{\zeta_{n}:n\geq 0\}$ are bounded ${\mathcal{F}}_{n}=\sigma(X_{0},\ldots,X_{n})$ adapted random variables. Let $U_{0}=0$ and for $n\geq 0$, let
@@
Suppose $\beta_{n}=1/(n+1)$, $t=1$ and $\zeta_{n}=1$ for all $n\geq 0$; then $U_{n}$ is just the average of $W_{1}$ through $W_{n}$. Therefore, if each $W_{n}$ has zero mean, then the almost sure convergence of $U_{n}$ to 0 is equivalent to $\{W_{n}:n\geq 1\}$ satisfying the strong law of large numbers (SLLN). It can be seen that if $\{W_{n}:n\geq 0\}$ are i.i.d. with mean 0 and finite variance, then (7) must hold in order for $U_{n}$ to converge to 00.
@@ 2.3 Stochastic Gradient Descent - Main Result
[…]
Now we define the iterative procedure. Choose some $y_{0}\in{\mathbb{R}}^{d}$, and define $Y_{0}=y_{0}$ For $n\geq 0$, let $\xi_{n}=\lambda_{n}(Y_{0},\ldots,Y_{n})$ and update $Y_{n}$ as folows:
$Y_{n+1}=Y_{n}-\sum_{i}\psi_{n,i}\eta_{n}\Bigl{(}\frac{\bigl{(}F(Y_{n}+e_{i}c_{n})+\xi_{n}M^{\prime}_{n+1,i}\bigr{)}-\bigl{(}F(Y_{n}-e_{i}c_{n})+\xi_{n}M^{\prime\prime}_{n+1,i}\bigr{)}}{2c_{n}}\Bigr{)}e_{i}.$ (15)
Here, as described in Section 1.1, $c_{n}$ is the increment, while $\eta_{n}$ is the step size.
@@
    The sequences $\{\beta_{n}:n\geq 0\}\subseteq(0,1)$ and $\{c_{n}:n\geq 0\}\subseteq(0,1)$ satify
@@
In the rest of the paper, the phrase “for all $n,i$” will mean “for all integers $n\geq 0$ and $i\in\{1,2,\ldots,d\}$.”
@@ 2.4 Sufficient Conditions for GSLLN
An $(F_{n})$-adapted sequence $\{W_{n}\}_{n\geq 0}$ satisfies the GSLLN at the rate $\{\beta_{n}\}$ if the following properties hold: for each $n\geq 1$, $i\in\{1,2,\ldots,d\}$, there exist sets $A_{n,i}\in{\mathcal{F}}_{n}$ such that, with
@@
An $(F_{n})$-adapted sequence $\{W_{n}\}_{n\geq m0}$ satisfies the GSLLN at the rate $\{\beta_{n}\}$ if the following properties hold: for each $n\geq 1$, $i\in\{1,2,\ldots,d\}$, there exist sets $A_{n,i}\in{\mathcal{F}}_{n}$ such that (20) and (22) hold; further, with $V_{n+1,i}$ defined by (19), we have
@@
Suppose $\{\beta_{n}:n\geq 0\}$ and $\{W_{n}:n\geq 1\}$ satisfy any one of the conditions $(H1)$$(H5)$ given below, in addition to $\{\beta_{n}:n\geq 0\}$ satisfying (7). Then $\{W_{n}:n\geq 1\}$ satisfies Generalized Strong Law of Large Numbers at the rate $\{\beta_{n}:n\geq 0\}$.
@@
    $\{W_{n}:n\geq 1\}$ are i.i.d. with $E(\|W_{1}\|^{\alpha})<\infty$ for some $21\leq\alpha<$ 2, $E(W_{1})=0$, and there exists a finite constant $D$ such that
@@
    This hypothesis consists of several conditions taken together. (i) $\{W_{n}:n\geq 1\}$ are i.i.d. with a symmetric distribution (that is, $W_{1}$ and $-W_{1}$ have the same distribtuion). (ii) There exists a constant $0<\delta\leq 1$ such that
@@
Suppose that one of the conditions $(K1)$$(K5)$ given below holds. Then GSLLN holds for $\{W_{n+1}:n\geq 0\}$ at the rate $\{\eta_{n}:n\geq 0\}$, and as a consequence, Stochastic Gradient Descent Theorem 5 holds (if in addition, (17), (13), (16), (14) hold).
@@
    Suppose $\{M^{\prime}_{n},M^{\prime\prime}_{n}:n\geq 1\}$ are i.i.d., for some $\alpha$, $21\leq\alpha<$ 2, $E(\|M^{\prime}_{1}\|^{\alpha})<\infty$, and for some finite constant D, $\beta_{0}\leq D$ and
@@ 3 Proofs
$\sum_{n}\tau_{n}=\infty,\;\;\;\sum_{n}\tau_{n}^{2}<\infty.$ (36)
@@
[$\bigl{[}\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr{]}+\sum_{k=m}^{n}\bigl{[}\prod_{\{k<j\leq n\}}(1-\tau_{j})\bigr{]}\tau_{k}=1,$ (38)
[$\bigl{[}\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr{]}s_{m}+\sum_{k=m}^{n}\bigl{[}\prod_{\{k<j\leq n\}}(1-\tau_{j})\bigr{]}\tau_{k}z_{k+1}={s_{n+1}}.$ (39)
@@
$\|s_{n+1}\|$ [$\leq\bigl{[}\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr{]}\|s_{m}\|+\sum_{k=m}^{n}\bigl{[}\prod_{\{k<j\leq n\}}(1-\tau_{j})\bigr{]}\tau_{k}\epsilon$
[$\leq\bigl{[}\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr{]}\|s_{m}\|+\epsilon.$
Taking limit as $n\rightarrow\infty$, we conclude that $\lim_{n}\|s_{n+1}\|\leq\epsilon$. Since this holds for all $\epsilon>0$, we conclude that $\|s_{n}\|\rightarrow 0$ completing proof of $(i)$. Coming to the proof of (ii), for $0\leq m\leq n$, let
@@
$\begin{split}{s_{n+1}}=&\bigl{[}\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr{]}s_{m}+\sum_{k=m}^{n}\bigl{[}\prod_{\{k<j\leq n\}}(1-\tau_{j})\bigr{]}\tau_{k}z_{k+1}.\\ =&\bigl{[}\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr{]}s_{m}+{\sum_{k=m}^{n}}[\prod_{\{k<j\leq n\}}(1-\tau_{j})](q_{k,n}-q_{k+1,n})\\ =&\bigl{[}\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr{]}s_{m}+[\prod_{\{m<j\leq n\}}(1-\tau_{j})]q_{m,n}\\ &+{\sum_{t=m+1}^{n}}[\prod_{\{t<j\leq n\}}(1-\tau_{j})-\prod_{\{t-1<j\leq n\}}(1-\tau_{j})]q_{t,n}\\ \end{split}$
@@
We will need another observation, which is needed to consider the case of i.i.d. errors when the mean may not be defined.
@@
In view of condition (20) and the Borel-Cantelli lemma, along with the fact that $\zeta_{n}$ is bounded, we can conclude that $\zeta_{n}U_{n+1,i}$ converges to zero almost surely. Hence in view of Lemma 10 it follows that $H_{n+1,i}$ converge to zero almost surely. If $V_{n+1,i}$ satisfies (24) then so does $\zeta_{n}V_{n+1,i}$. As a consequence, $J_{n+1,i}$ converges to 0 in view of (i) in Lemma 10 while if $V_{n+1,i}$ satisfies (21) then again so does $\zeta_{n}V_{n+1,i}$. As a consequence $J_{n+1,i}$ converges to 0 in view of (ii) in Lemma 10. It is clear that
@@
$\sum_{k=0}^{n}\beta_{k}\zeta_{n}R_{k+1,i}\;\text{ converges almost surely.}$. (46)
@@
We are going to show that for every $\omega\in\Omega$ such that $S_{n+1}(\omega)$ defined by (49) converges, $X_{n+1}(\omega)-x^{*}$ defined by (47) converges to zero. We will drop $\omega$ in the notation as usual, and we should keep in mind that the choices of $n,\,m,\,K$ in what follows can depend upon $\omega$.
@@
[…]
@@
$\begin{split}L_{m,n}=&{\sum_{k=m}^{n}}\phi_{k}[\prod_{\{k<j\leq n\}}(1-\tau_{j})]\tau_{k}Z_{k+1}\\ =&\phi_{m}{\sum_{k=m}^{n}}[\prod_{\{k<j\leq n\}}(1-\tau_{j})]\tau_{k}Z_{k+1}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+{\sum_{k=m+1}^{n}}(\phi_{k}-\phi_{m})[\prod_{\{k<j\leq n\}}(1-\tau_{j})]\tau_{k}Z_{k+1}\\ =&\phi_{m}D_{m,n}+{\sum_{k=m+1}^{n}}(\phi_{k}-\phi_{m})[\prod_{\{k<j\leq n\}}(1-\tau_{j})]\tau_{k}Z_{k+1}\\ \end{split}$ (56)
@@
First, using $\rho<1$, get $\epsilon>0$ such that $\epsilon<\frac{1}{3}$
@@
$\begin{split}\|V_{m,n}\|&\leq\sum_{k=m}^{n}\bigl{[}\prod_{{}_{\{j:k<j\leq n\}}}(1-\tau_{j})\bigr{]}\tau_{k}\|((X_{k}-x^{*})-R_{k})\|\\ &\leq\sum_{k=m}^{n}\bigl{[}\prod_{{}_{\{j:k<j\leq n\}}}(1-\tau_{j})\bigr{]}\tau_{k}\rho\|(X_{k}-x^{*})\|\\ &\leq\rho\sum_{k=m}^{n}\bigl{[}\prod_{{}_{\{j:k<j\leq n\}}}(1-\tau_{j})\bigr{]}\tau_{k}\theta_{k}\\ &\leq\rho\theta_{m}\\ \end{split}$ (72)
@@
(Of Theorem 5. Several steps in the proof are the same as in the proof of Theorem 4. Recall that $\xi_{n}=\lambda_{n}(Y_{0},Y_{1},\ldots,Y_{n})$ where $\lambda_{n}$ satisfies (13) and let $\phi_{n}=(1+\max_{0\leq k\leq n}\|Y_{k}\|_{\infty})$. Let $b$ be such that (17) holds and let $\zeta_{n}=-b^{-1}\phi_{n}^{-1}\xi_{n}$, $T_{n,i}=\frac{1}{2bc_{n}}\bigl{(}F(Y_{n}+e_{i}c_{n})-F(Y_{n}-e_{i}c_{n}))$, $Z_{n+1,i}=\psi_{n,i}\zeta_{n}W_{n+1,i}$, $\tau_{n}=b\eta_{n}$. Then $\tau_{n}\phi_{n}Z_{n+1,i}=b\eta_{n}\psi_{n,i}\phi_{n}\zeta_{n}W_{n+1,i}=-\eta_{n}\psi_{n,i}\xi_{n}M_{n+1,i}$. Note that $\{\zeta_{n}:n\geq 0\}$ is bounded. For later use, let us note that in view of assumption (17) on $F$, it follows that
@@
Recall that $\{\eta_{n},c_{n},\psi_{n,i}:n\geq 0,1\leq i\leq d\}$ satisfy (16), (14). Thus $\sum_{n}\psi_{n,i}\tau_{n}=\infty$ almost surely. We are going to show that for every $\omega\in\Omega$ such that $S_{n+1,i}(\omega)$ defined by (77) converges and $\sum_{n}\psi_{n,i}(\omega)\tau_{n}=\infty$ for each $i$; $Y_{n+1,i}(\omega)-x^{*}_{i}$ defined by (76) converges to zero, as in proof of Theorem 4. We will drop $\omega$ in the notation as usual, and we should keep in mind that choices of $n,\,m,\,K$ in what follows can depend upon $\omega$.
@@
[…]
@@
$\begin{split}L_{m,n,i}=&{\sum_{k=m}^{n}}\phi_{k}[\prod_{\{k<j\leq n\}}(1-\psi_{j,i}\tau_{j})]\psi_{k,i}\tau_{k}Z_{k+1,i}\\ =&\phi_{m}{\sum_{k=m}^{n}}[\prod_{\{k<j\leq n\}}(1-\psi_{j,i}\tau_{j})]\psi_{k,i}\tau_{k}Z_{k+1,i}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+{\sum_{k=m+1}^{n}}(\phi_{k}-\phi_{m})[\prod_{\{k<j\leq n\}}(1-\psi_{j,i}\tau_{j})]\psi_{k,i}\tau_{k}Z_{k+1,i}\\ =&\phi_{m}D_{m,n}+{\sum_{k=m+1}^{n}}(\phi_{k}-\phi_{m})[\prod_{\{k<j\leq n\}}(1-\psi_{j,i}\tau_{j})]\psi_{k,i}\tau_{k}Z_{k+1,i}\\ \end{split}$ (88)
@@
$\begin{split}|V_{m,n,i}|&\leq\sum_{k=m}^{n}\bigl{[}\prod_{{}_{\{j:k<j\leq n\}}}(1-\psi_{j,i}\tau_{j})\bigr{]}\psi_{k,i}\tau_{k}\rho|Y_{k,i}-x^{*}_{i}|\\ &\leq\sum_{k=m}^{n}\bigl{[}\prod_{{}_{\{j:k<j\leq n\}}}(1-\psi_{j,i}\tau_{j})\bigr{]}\tau_{k}\rho\theta_{m}\\ &\leq\rho\theta_{m}\\ \end{split}$ (106)

### v3

@@ 1.1 Historical Review
$\hat{G}(Y_{n})=\frac{1}{2c_{n}}[(F(Y_{n}+c_{n})+M^{\prime}_{n+1})-(F(Y_{n}-c_{n})+M^{\prime\prime}_{n+1})],$ (3)
@@
Subsequently, Blum [3, 4] extended the approach to higher dimensions, and also proved that the iterations converge almost surely, provided (5) hold. For this reason, the conditions (5) are referred to as the Kiefer-Wolfowitz-Blum (KWB) conditions.
@@ 1.2 Contributions of the Paper
    In the special case where $\{W_{n}\}$ are i.i.d. with a symmetric distribution (that is, $W_{1}$ and $W_{1}$ have the same distribution), GSLLN holds if there exists a $\delta\in(0,1]$ such that
@@ 2.2 Stochastic Approximation - Main Result
Let $d\geq 1$ be a fixed integer and $G:{\mathbb{R}}^{d}\mapsto{\mathbb{R}}^{d}$ be a function with a unique zero, call it $x^{*}$. Thus $G(x^{*})=0$ and $G(x)\neq 0$ for all $x\neq x^{*}$. We consider a generalization of (1), in which the noise $W_{n+1}$ is multiplied by a coefficient that can depend on the past history of the iterations. The set up is as follows: For $n\geq 0$, let $\lambda_{n}:({\mathbb{R}}^{d})^{n+1}\mapsto{\mathbb{R}}$ be such that, for some constant $C_{1}$ and some norm $\|\cdot\|$ on ${\mathbb{R}}^{d}$, $\lambda_{n}$ satisfies
@@
Note that, while $\|\cdot\|$ can be anything, it is essential to use the maximum of $\|u_{0}\|,\cdots,\|u_{n}\|$ in (9). This avoids the norm becoming larger and larger as $n$ increases. Now the iterative scheme is defined as follows: For $x_{0}\in{\mathbb{R}}^{d}$ fixed, $X_{0}=x_{0}$ and for $n\geq 0$ let
@@
Our aim is to establish conditions on the step size sequence $\{\beta_{n}:n\geq 0\}$ and noise sequence $\{W_{n}:n\geq 1\}$ that ensure the almost sure convergence of $\{X_{n}:n\geq 0\}$ defined by (10) to $x^{*}$. Though all norms on ${\mathbb{R}}^{d}$ are equivalent, the fact that (11) holds for one particular does not imply that (11) holds for some other norm. Thus the assumption is that there exists some suitable norm $\|\cdot\|$ on ${\mathbb{R}}^{d}$ such that (11) holds, and it does not matter which norm it is. Note that the constants $b$ and $\rho$ appearing in (11) are used only in the proofs, and do not appear in the iteration formula.
@@
Suppose $H:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ is a contraction with respect to some norm $\|\cdot\|$. Thus there exists a constant $\rho\in[0,1)$ such that
@@
Let $x^{*}$ denote the unique fixed point of $H$, and define G(x)=xH(x)G(x)=x-H(x). Then clearly $x^{*}$ is the unique solution of $G(x)={\bf 0}$. Moreover, it is easy to see that (11) is satisfied with $b=1$. This is because
@@
where for symmetric matrices $A$ and $B$, $A\leq B$ is equivalent to BAB-A being positive semidefinite. The above condition is also equivalent to the statement that the largest eigenvalue of $M(x)$ is bounded above by $h<\infty$, and the smallest eigenvalue of $M(x)$ is bounded below by $g>0$. With these assumptions, it is shown that (7) holds with $\|\cdot\|=\|\cdot\|_{2}$, and
$b=\frac{1}{1+g+h},\quad\rho=\frac{1+h}{1+g+h}<1.$
@@
$G(x)$ $=\int_{0}^{1}\frac{d}{d\lambda}G(x^{*}+\lambda(x-x^{*}))\;d\lambda$
$=\int_{0}^{1}M(x^{*}+\lambda(x-x^{*}))\cdot(x-x^{*})\;d\lambda.$
@@
$x-x^{*}-bG(x)=\int_{0}^{1}[I-bM(x^{*}+\lambda(x-x^{*}))]\cdot(x-x^{*})\;d\lambda.$
By assumption, the spectrum of the symmetric matrix $M(x^{*}+\lambda(x-x^{*}))$ lies in the interval $[g,h]$. Hence the spectrum of $I-bM(x^{*}+\lambda(x-x^{*}))$ lies in the interval $1-\frac{h}{1+g+h},1-\frac{g}{1+g+h}]$. Since both numbers are positive, the spectral norm (which is the same as the $\ell_{2}$-induced norm) satisfies
$\|I-bM(x^{*}+\lambda(x-x^{*}))\|_{2\rightarrow 2}\leq 1-\frac{g}{1+g+h}=\rho.$
@@
$\|x-x^{*}-bG(x)\|\leq\int_{0}^{1}\rho\|x-x^{*}\|_{2}\;d\lambda=\rho\|x-x^{*}\|_{2}.$
To motivate the concept of the GSLLN, let us study a special case, when $G(x)=tx$, $t>0$ and $\{\zeta_{n}:n\geq 0\}$ are bounded ${\mathcal{F}}_{n}=\sigma(X_{0},\ldots,X_{n})$ adapted random variables. Let $U_{0}=0$ and for $n\geq 0$, let
@@
Suppose $\beta_{n}=1/(n+1)$, $t=1$ and $\zeta_{n}=1$ for all $n\geq 0$; then $U_{n}$ is just the average of $W_{1}$ through $W_{n}$. Therefore, if each $W_{n}$ has zero mean, then the almost sure convergence of $U_{n}$ to 0 is equivalent to $\{W_{n}:n\geq 1\}$ satisfying the strong law of large numbers (SLLN). It can be seen that if $\{W_{n}:n\geq 0\}$ are i.i.d. with mean 0 and finite variance, then (7) must hold in order for $U_{n}$ to converge to 0.
@@ 2.3 Stochastic Gradient Descent - Main Result
Let $\{M^{\prime}_{m},M^{\prime\prime}_{m}:m\geq 1\}$ be sequences of ${\mathbb{R}}^{d}$ valued random variables, representing observational errors appearing in (15) below). Let $\{\psi_{n,i}:n\geq 0,\,1\leq i\leq d\}$ be $\{0,1\}$ valued $({\mathcal{F}}^{M^{\prime},M^{\prime\prime}})$ adapted random variables. We will also assume that $\forall i\in\{1,2,\ldots,d\}$
$\sum_{n=0}^{\infty}\psi_{n,i}\eta_{n}=\infty\mbox{ a.s.}$ (14)
where, as described in Section 1.1, $\eta_{n}$ is the step size and $c_{n}$ that appears below is the increment. Now we define the iterative procedure. Choose some $y_{0}\in{\mathbb{R}}^{d}$, and define $Y_{0}=y_{0}$ . For $n\geq 0$, let $\xi_{n}=\lambda_{n}(Y_{0},\ldots,Y_{n})$ and update $Y_{n}$ as folows:
$Y_{n+1}=Y_{n}-\sum_{i}\psi_{n,i}\eta_{n}\Bigl(\frac{\bigl(F(Y_{n}+e_{i}c_{n})+\xi_{n}M^{\prime}_{n+1,i}\bigr)-\bigl(F(Y_{n}-e_{i}c_{n})+\xi_{n}M^{\prime\prime}_{n+1,i}\bigr)}{2c_{n}}\Bigr)e_{i}$ (15)
where $e_{i}$ is the unit vector in the $i^{th}$ direction.
@@
    The sequences $\{\eta_{n}:n\geq 0\}\subseteq(0,1)$ and $\{c_{n}:n\geq 0\}\subseteq(0,1)$ satisfy
@@
In the rest of the paper, the phrase “for all n,in,i” will mean “for all integers $n\geq 0$ and $i\in\{1,2,\ldots,d\}$.”
@@ 2.4 Sufficient Conditions for GSLLN
An $(F_{n})$-adapted sequence $\{W_{n}\}_{n\geq 0}$ satisfies the GSLLN at the rate $\{\beta_{n}\}$ w.r.t. the filtration $({\mathcal{F}}_{n})_{n\geq 0}$ if the following properties hold: for each $n\geq 1$, $i\in\{1,2,\ldots,d\}$, there exist sets $A_{n,i}\in{\mathcal{F}}_{n}$ such that, with
@@
An $(F_{n})$-adapted sequence $\{W_{n}\}_{n\geq 0}$ satisfies the GSLLN at the rate $\{\beta_{n}\}$ w.r.t. the filtration $({\mathcal{F}}_{n})_{n\geq 0}$ if the following properties hold: for each $n\geq 1$, $i\in\{1,2,\ldots,d\}$, there exist sets $A_{n,i}\in{\mathcal{F}}_{n}$ such that (20) and (22) hold; further, with $V_{n+1,i}$ defined by (19), we have
@@
Suppose $\{\beta_{n}:n\geq 0\}$ and $\{W_{n}:n\geq 1\}$ satisfy any one of the conditions (H1)(H1)$(H5)$ given below, in addition to $\{\beta_{n}:n\geq 0\}$ satisfying (7). Then $\{W_{n}:n\geq 1\}$ satisfies Generalized Strong Law of Large Numbers at the rate $\{\beta_{n}:n\geq 0\}$ w.r.t. the filtration $({\mathcal{F}}^{W}_{n})_{n\geq 0}$.
@@
    $\{W_{n}:n\geq 1\}$ are i.i.d. with $E(\|W_{1}\|^{\alpha})<\infty$ for some $1\leq\alpha<2$, $E(W_{1})=0$, and there exists a finite constant $D$ such that
@@
    This hypothesis consists of several conditions taken together. (i) $\{W_{n}:n\geq 1\}$ are i.i.d. with a symmetric distribution (that is, $W_{1}$ and $W_{1}$ have the same distribtuion). (ii) There exists a constant $0<\delta\leq 1$ such that
@@
Suppose that one of the conditions (K1)(K1)$(K5)$ given below holds. Then GSLLN holds for $\{W_{n+1}:n\geq 0\}$ at the rate $\{\eta_{n}:n\geq 0\}$ w.r.t. $({\mathcal{F}}^{M^{\prime},M^{\prime\prime}})$, and as a consequence, Stochastic Gradient Descent Theorem 5 holds (if in addition, (17), (13), (16), (14) hold).
@@
    Suppose $\{M^{\prime}_{n},M^{\prime\prime}_{n}:n\geq 1\}$ are i.i.d., for some $\alpha$, $1\leq\alpha<2$, $E(\|M^{\prime}_{1}\|^{\alpha})<\infty$, and for some finite constant D, $\beta_{0}\leq D$ and
@@ 3 Proofs
$\lim_{n\rightarrow\infty}\tau_{n}=0,\;\;\;\sum_{n=0}^{\infty}\tau_{n}=\infty.$ (36)
@@
$\bigl[\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr]+\sum_{k=m}^{n}\bigl[\prod_{\{k<j\leq n\}}(1-\tau_{j})\bigr]\tau_{k}=1,$ (38)
$\bigl[\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr]s_{m}+\sum_{k=m}^{n}\bigl[\prod_{\{k<j\leq n\}}(1-\tau_{j})\bigr]\tau_{k}z_{k+1}={s_{n+1}}.$ (39)
@@
$\|s_{n+1}\|$ $\leq\bigl[\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr]\|s_{m}\|+\sum_{k=m}^{n}\bigl[\prod_{\{k<j\leq n\}}(1-\tau_{j})\bigr]\tau_{k}\epsilon$
$\leq\bigl[\prod_{\{m\leq j\leq n\}}(1-\tau_{j})\bigr]\|s_{m}\|+\epsilon.$
Taking limit as $n\rightarrow\infty$, we conclude that $\lim_{n}\|s_{n+1}\|\leq\epsilon$. Since this holds for all $\epsilon>0$, we conclude that $\|s_{n}\|\rightarrow 0$ completing proof of $(i)$.
Coming to the proof of (ii), for $0\leq m\leq n$, let
@@
[…]
@@
We will need another observation to consider the case of i.i.d. errors when the mean may not be defined.
@@
In view of condition (20) and the Borel-Cantelli lemma we can conclude that $\zeta_{n}U_{n+1,i}$ converges to zero almost surely. Hence in view of Lemma 10 it follows that $H_{n+1,i}$ converge to zero almost surely. If $V_{n+1,i}$ satisfies (24) then so does $\zeta_{n}V_{n+1,i}$, because $\{\zeta_{n}\}$ is bounded. As a consequence, $J_{n+1,i}$ converges to 0 in view of (i) in Lemma 10 while if $V_{n+1,i}$ satisfies (21) then again so does $\zeta_{n}V_{n+1,i}$. As a consequence $J_{n+1,i}$ converges to 0 in view of (ii) in Lemma 10. It is clear that
@@
$\sum_{k=0}^{n}\beta_{k}\zeta_{n}R_{k+1,i}\;\text{ converges almost surely.}$ (46)
@@
We are going to show that for every $\omega\in\Omega$ such that $S_{n+1}(\omega)$ defined by (49) converges, $X_{n+1}(\omega)-x^{*}$ defined by (47) converges to zero. We will drop $\omega$ in the notation as usual, and we should keep in mind that the choices of n,m,Kn,\,m,\,K in what follows can depend upon $\omega$.
@@
Let $\gamma_{n}=\max\{\|X_{k}-x^{*}\|,\;\;0\leq k\leq n\}$ and $\gamma^{*}=\sup_{n}\gamma_{n}$. Then $\|X_{k}\|\leq(\gamma_{n}+\|x^{*}\|)$ for $0\leq k\leq n$, and recalling that $\phi_{n}=(1+\max_{0\leq k\leq n}\|X_{k}\|)$, it follows that
@@
[…]
@@
First, using $\rho<1$, get $\epsilon>0$ such that $\epsilon<\frac{1}{3}$ and
@@
[…]
@@
Of Theorem 5. Several steps in the proof are the same as in the proof of Theorem 4. Recall that $\xi_{n}=\lambda_{n}(Y_{0},Y_{1},\ldots,Y_{n})$ where $\lambda_{n}$ satisfies (13) and let $\phi_{n}=(1+\max_{0\leq k\leq n}\|Y_{k}\|_{\infty})$. Let $b$ be such that (17) holds and let $\zeta_{n}=-b^{-1}\phi_{n}^{-1}\xi_{n}$, $T_{n,i}=\frac{1}{2bc_{n}}\bigl(F(Y_{n}+e_{i}c_{n})-F(Y_{n}-e_{i}c_{n}))$, $Z_{n+1,i}=\psi_{n,i}\zeta_{n}W_{n+1,i}$, $\tau_{n}=b\eta_{n}$. Then $\tau_{n}\phi_{n}Z_{n+1,i}=b\eta_{n}\psi_{n,i}\phi_{n}\zeta_{n}W_{n+1,i}=-\eta_{n}\psi_{n,i}\xi_{n}M_{n+1,i}$. Note that $\{\zeta_{n}:n\geq 0\}$ is bounded. For later use, let us note that in view of assumption (17) on $F$, it follows that
@@
Recall that $\{\eta_{n},c_{n},\psi_{n,i}:n\geq 0,1\leq i\leq d\}$ satisfy (16), (14). Thus $\sum_{n}\psi_{n,i}\tau_{n}=\infty$ almost surely. We are going to show that for every $\omega\in\Omega$ such that $S_{n+1,i}(\omega)$ defined by (77) converges and $\sum_{n}\psi_{n,i}(\omega)\tau_{n}=\infty$ for each $i$; $Y_{n+1,i}(\omega)-x^{*}_{i}$ defined by (76) converges to zero, as in proof of Theorem 4. We will drop $\omega$ in the notation as usual, and we should keep in mind that choices of n,m,Kn,\,m,\,K in what follows can depend upon $\omega$.
@@
Let $\gamma_{n}=\max\{\|Y_{k}-x^{*}\|_{\infty}:\;0\leq k\leq n\}$ and $\gamma^{*}=\sup_{n}\gamma_{n}$. Note that for all $i$ and $k\leq n$, $|Y_{k,i}-x^{*}_{i}|\leq\gamma_{n}$. By definition, we have $\|Y_{n}\|_{\infty}\leq(\gamma_{n}+\|x^{*}\|_{\infty})$ and recalling that $\phi_{n}=(1+\max_{0\leq k\leq n}\|Y_{k}\|_{\infty})$, it follows that
@@
[…]
@@
[…]